The information in the chart below compares the price of several products.

chest of drawers entertainment center medicine cabinet side table sofa
$255 $440 $140 $90 $430

Write an equation or inequality using the appropriate symbol = , < , or > to compare the prices of the sofa and the chest of drawers. Write each price exactly as it appears in the table, including the dollar sign. For example $440>$90 .
Answer #1

SOLUTION

The price of several products are as following :

chest of drawers entertainment center medicine cabinet side table sofa
$255 $440 $140 $90 $430

We have to write an equation or inequality using the appropriate symbol =, <, or > to compare the prices of the sofa and the chest of drawers.

Relation

Relation Between Price of Sofa and Price of the chest of drawers is

Let A * (Price of Sofa) = (Price of the chest of drawers)

\small \RightarrowA * [$430 ]= [$255]

\small \Rightarrow A = [$255] / [$430 ]

\small \Rightarrow A= 51/86

I.e (51/86) * Price of Sofa = (Price of the chest of drawers)

Comparison

Price of Sofa > Price of the chest of drawers

i.e $430 > $255